OpenStudy (anonymous):

Need help on my homework. I've done a problem like this before but with a simpler expression in the parenthesis. I am not sure how to even approach the problem. The question is: Draw a right triangle to simplify the given expression. Assume x > 0. sin(2cos^1x). (Hint: use sin2(theta) = 2sin(theta)cos(theta).

zepdrix (zepdrix):

Was that meant to be cos^(-1)x inside? So the inverse function of cosine? arccos(x)

OpenStudy (anonymous):

Yes. Sorry for the confusion

zepdrix (zepdrix):

Remember that the trig functions do something kind of interesting. They take an angle... and then the trig "function" applied to this angle gives a relationship based on sides of a triangle. \[\large\rm \cos \theta=\frac{adjacent}{hyptenuse}\]

zepdrix (zepdrix):

Inverse cosine does the opposite, it takes a value or ratio of sides, and spits out an angle.\[\large\rm \color{orangered}{\cos^{-1}x=\theta}\]

OpenStudy (anonymous):

It is actually 2cos^(-1)x inside the parentheses. Is there some way to factor out the 2?

zepdrix (zepdrix):

We'll worry about that later :)

zepdrix (zepdrix):

We would like to do some triangle business first.

zepdrix (zepdrix):

So we're saying that our inverse cosine function is equivalent to some angle theta,\[\large\rm \color{orangered}{\cos^{-1}x=\theta}\]Do you understand how to get from there, \[\large\rm \cos \theta=x\]to here?

OpenStudy (anonymous):

Yes I just worked it out and rearranged it

zepdrix (zepdrix):

So anyway, this cos(theta)=x we can write as,\[\large\rm \cos \theta=\frac{x}{1}=\frac{adjacent}{hypotenuse}\]And we can draw a right triangle to illustrate this relationship.

zepdrix (zepdrix):